3 / [ 2x - 3] = 5x+7 / [x + 1] Solve using quadratic formula The answer should be -1.4 and 1.8?
This is a fraction question. I'm not sure if I used parentheses correct

Answer by stanbon(75887)

3 / [ 2x - 3] = 5x+7 / [x + 1] Solve using quadratic formula The answer should be -1.4 and 1.8?
----
Cross-multiply to get:
3(x+1) = (2x-3)(5x+7)
----
3x+3 = 10x^2 -x - 21
----
10x^2 - 4x - 24 = 0
----
Simplify::
5x^2 - 2x - 12 = 0
---
Quad Formula::
x = [2 +- sqrt(4-4*5*-12)]/10
------
x = [2 +- sqrt(244)]/10
------
x = [2 +- 2sqrt(61)]/10
----
Answers are not what you posted.
Cheers,
Stan H.
----------